(5^7)^2 ÷ (5^3)^-6 x (5^9)^-4
I got an answer of 1/5^68 and then =1/3.38813
not sure if i am correct. i have the work too.

Answer by Alan3354(69443)

(5^7)^2 ÷ (5^3)^-6 x (5^9)^-4
I got an answer of 1/5^68 and then =1/3.38813
------------------
= (5^14/(5^-18)) * 5^-36 Maybe, it's not clear how much is in the DEN.
= 5^32 * 5^-36
= 5^-4
= 0.0016
------------
Use more parenthese to clarify what goes where.
----------
1/5^68 is not 1/3.38813